# Importance Ranking in Complex Networks via Influence-aware Causal Node Embedding

### v2

@@ 5.1 Description of datasets
We train the model on five synthetic graphs and evaluate it on six real-world networks to validate its effectiveness.
Training datasets. Five representative synthetic network models are used for training, including Barabási–Albert (BA) [barabasi2009scale], Extreme Homogeneous (EH) [lou2020towards], Erdös-Rényi (ER) random-graph [erdHos1961strength], Q-Snapback (QS) [lou2018toward] and Random Hexagon (RH) [lou2018toward]. The BA networks are generated using the preferential attachment mechanism, where new nodes are sequentially added and connected to existing nodes with a probability proportional to their current degrees, leading to a scale-free degree distribution. The EH networks are obtained by performing additional random edge rectifications on ER networks to transform their degree distribution from Poisson to near-uniform. The ER random graphs are generated by connecting each pair of nodes independently with a fixed probability $p$, producing networks with a binomial degree distribution. The QS networks consist of a directed backbone chain with multiple probabilistic “snapback” edges linking newly added nodes to previously added ones within a defined range, thereby enhancing backward connectivity and robustness. The RH networks are composed of randomly connected hexagonal substructures that capture local clustering and spatial organization similar to lattice-like systems. All trained networks comprise 1,000 nodes with an average degree of 4.
@@
  • Karate [zachary1977information]: A human social network, in which each node represents a member in the club and each link shows the relationship between two members.
  • Jazz [gleiser2003community]: A human collaboration network, in which each node represents a jazz musician and each link indicates that two musicians have performed together in the same band.
  • Email-univ [guimera2003self]: An email communication network from the University Rovira i Virgili in Spain, where nodes represent users and edges indicate at least one email was exchanged.
  • USAir [kunegis2013konect]: A directed weighted network represents the flight connections among 1,574 U.S. airports in 2010. Each node corresponds to an airport, and each directed edge indicates the existence of at least one flight from the origin to the destination airport. The edge weight quantifies the total number of flights operated on that route during the year.
  • Vidal [rual2005towards]: A network represents an initial version of a proteomescale map of Human binary protein–protein interactions.
  • Email-dnc [rossi2015network]: A directed unweighted network derived from the 2016 Democratic National Committee (DNC) email leak. Each node represents an individual user, and each directed edge indicates that at least one email was sent from one user to another.
@@ 5.2 Comparative methods
  • Degree Centrality (DC) [bonacich1972factoring]. It quantifies the influence of a node based on the number of direct connections it has to other nodes in the graph. The degree centrality of a node u is formally defined as
@@
  • Betweenness Centrality (BC) [freeman1977set]. It measures the importance of a node by quantifying how frequently it appears on the shortest paths between all pairs of nodes in the network.
@@
  • Eigenvector Centrality (EC) [bonacich1987power]. It assigns each node a centrality score based not only on its own connectivity but also on the importance of the nodes it is connected to. The EC score $\bm{X}_{u}$ of node $u$ satisfies
@@
  • $H$-index (HI) [hirsch2005index]. The $H$-index of a node considers the degrees of its neighbors and reflects how many of them are themselves influential. For a node $u$ with neighbors $j_{1},j_{2},\ldots,j_{k}$, the $H$-index is defined as
@@
  • $K$-Shell (KS) [kitsak2010identification]. The $K$-shell algorithm operates through an iterative pruning process. It iteratively peels away network layers to assign each node a $k$-value indicating its coreness based on residual connections. A higher $K$-shell value indicates greater node influence.
  • GNN-Bet [maurya2021graph]. This method enables node ranking by employing a dual-path aggregation mechanism that separately propagates features along incoming and outgoing shortest paths using row-wise modified adjacency matrices, with final ranking scores generated through multiplicative combination of the aggregated path representations.
  • GNN-Close [maurya2021graph]. This approach achieves node ranking through a hierarchical feature aggregation scheme that combines normal adjacency operations in the initial layer with column-wise modified matrices in subsequent layers to constrain feature propagation along shortest paths, producing final ranking scores via summation of layer-wise outputs.
  • RCNN [yu2020identifying]. This method leverages convolutional neural networks to extract features from node-centric subgraph structures and ranks node importance based on learned representations.
  • CGNN [zhang2022new]. This method enhances RCNN by incorporating additional GNN layers, improving the capture of topological dependencies for more accurate vital node identification.
  • AGNN [xiong2024vital]. This method combines a GCN-based autoencoder with a GNN ranking head to jointly learn structural embeddings and predict node importance using listwise ranking loss.
  • STGCN [huang2025sparse]  [^1] Since AGNN and STGCN employ the adjacency matrix as the input to their autoencoder and do not explicitly address the issue of inconsistent numbers of nodes between training and testing networks, we instead incorporate the generated node features by the adjacency matrix as the input to the AGNN and STGCN models.. This method leverages a hybrid architecture combining Sparse Transformer layers for global dependency modeling and GCNs for local topological features to rank key nodes in complex networks.
@@ 5.3 Evaluation criterion
In this section, we introduce the evaluation metric used to measure the effectiveness of the predicted ranking results by ICAN, Kendall’s $\tau$ coefficient [knight1966computer]. This non-parametric statistic is widely employed to measure the ordinal association between two ranking lists. It quantifies the similarity between two ranking orders by comparing the numbers of concordant and discordant pairs. The definition of $\tau$ is
@@
where $n$ is the number of items (or nodes), $N_{c}$ is the number of concordant pairs, $N_{d}$ is the number of discordant pairs, and $\tfrac{1}{2}n(n-1)$ is the total number of possible item pairs. For two rankings $\{(x_{i},y_{i})\}$ and $\{(x_{j},y_{j})\}$, a pair $(i,j)$ is considered concordant if $(x_{i}-x_{j})(y_{i}-y_{j})>0$, discordant if $(x_{i}-x_{j})(y_{i}-y_{j})<0$, and neither if $x_{i}=x_{j}$ or $y_{i}=y_{j}$. The value of Kendall’s $\tau$ lies between 1-1 and $1$: a value of $1$ indicates complete agreement between rankings, 1-1 indicates complete disagreement, and 0 indicates no correlation. This metric is particularly useful for evaluating graph-based ranking models, as it reflects how well the predicted node ranking preserves the ground-truth ordering.
@@ 5.4 Implementation details
The parameter settings for ICAN and the comparative methods are as follows.
For the ICAN method, set the number of hidden layers to be $l=5$ and each hidden layer to have 32 neurons. Let
@@
The learning rate is $\mu=0.001$ and the maximum number of iterations $T=10$. In addition, we set the recovery probability $\delta=1$ and the infection probability $\gamma=1.5\times\gamma_{c}$ to ensure effective spreading within the network. The infection threshold $\gamma_{c}$ is derived from mean-field theory [moreno2002epidemic] as
@@ 5.5.1 Training on BA-generated network and testing on various real-world networks
In the first experiment, we conduct experiments by training on the BA generated network and testing on different real-world networks to validate the model’s effectiveness. Table II presents the predicted Kendall’s correlation coefficients across different datasets. The experimental results comprehensively evaluate the proposed ICAN model against several baseline methods for node ranking across six real-world networks. As we can see, ICAN demonstrates superior and consistent performance. Notably, ICAN achieves the highest Kendall’s coefficient on every individual dataset. This culminates in the best overall average Kendall’s coefficient of 0.7707, securing the top rank among all compared methods. The fact that ICAN, trained on a BA synthetic network, generalizes effectively to diverse real-world networks highlights its robust cross-network predictive capability. This can be attributed to the ICAN learning network-invariant node embeddings that transfer reliably across different networks.
@@
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Average Rank
DC 0.6749 0.7903 0.6236 0.6119 0.4648 0.5613 0.6211 6
BC 0.5685 0.4643 0.2374 0.4407 0.5876 0.4032 0.4503 12
EC 0.8069 0.8212 0.2534 0.6554 0.2452 0.5503 0.5554 8
HI 0.6654 0.8271 0.6854 0.6284 0.7568 0.5672 0.6884 4
KS 0.6421 0.7713 0.6828 0.6369 0.7619 0.4988 0.6656 5
GNN-Bet 0.6530 0.6789 0.3645 0.5350 0.2866 0.3662 0.4807 10
GNN-Close 0.5352 0.6551 0.3629 0.5818 0.2429 0.4442 0.4704 11
RCNN 0.7528 0.8455 0.7522 0.6709 0.9035 0.5432 0.7447 2
CGNN 0.7058 0.7911 0.6488 0.5835 0.9442 0.5649 0.7064 3
AGNN 0.6505 0.7584 0.3100 0.6429 0.4108 0.4946 0.5445 9
STGCN 0.5640 0.7000 0.7327 0.3972 0.5784 0.4248 0.5662 7
ICAN 0.8090 0.8504 0.7625 0.6758 0.9524 0.5740 0.7707 1
@@ 5.5.2 Training on ER-generated network and Testing on various real-world networks
Next, we conduct experiments by training on the ER generated network and testing on different real-world networks. Table III presents the Kendall’s $\tau$ coefficients of various methods evaluated across six real-world networks. As shown, ICAN attains the highest Kendall’s $\tau$ values on four out of six datasets, with an average coefficient of 0.7732, ranking first overall. This clearly demonstrates its strong generalization ability when trained on the ER-generated network and tested on structurally diverse real-world networks.
@@
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Average Rank
DC 0.6749 0.7903 0.6236 0.6119 0.4648 0.5613 0.6211 6
BC 0.5685 0.4643 0.2374 0.4407 0.5876 0.4032 0.4503 12
EC 0.8069 0.8212 0.2534 0.6554 0.2452 0.5503 0.5554 7
HI 0.6654 0.8271 0.6854 0.6284 0.7568 0.5672 0.6884 3
KS 0.6421 0.7713 0.6828 0.6369 0.7619 0.4988 0.6656 5
GNN-Bet 0.6137 0.6394 0.3744 0.5565 0.4125 0.4566 0.5089 10
GNN-Close 0.6887 0.2779 0.4550 0.5470 0.4079 0.4122 0.4648 11
RCNN 0.7094 0.8014 0.8141 0.6823 0.9119 0.5490 0.7447 2
CGNN 0.6749 0.6214 0.6323 0.5835 0.9442 0.5649 0.6702 4
AGNN 0.7622 0.4889 0.2950 0.6829 0.4008 0.4534 0.5139 9
STGCN 0.5568 0.6525 0.7141 0.3952 0.5792 0.3735 0.5452 8
ICAN 0.8092 0.8460 0.8228 0.6545 0.9524 0.5540 0.7732 1
@@ 5.6.2 Ablation study for the feature–task co-optimization mechanism
In this study, we integrate the causal ranking loss and causal reconstruction loss within a unified framework to enable causal feature extraction and ranking tasks. To evaluate the effectiveness of the proposed end-to-end feature–task co-optimization mechanism, we conduct an ablation study by transforming the integrated model into a two-stage implementation, in which the feature extraction and ranking processes are decoupled, thereby removing the synergistic optimization effect. This variant is referred to as ICAN_w/o_Syner. In the first stage, the causal reconstruction loss and a regularization term are employed as the loss function to optimize the causal node embedding module, yielding low-dimensional node representations that comprehensively encapsulate the topological properties of the network. During the second stage, the learned node embeddings are fed into the ranking prediction module, which is optimized using the causal ranking loss along with a regularization term to generate the predicted node influence score. These predictions are then compared against the node influence score derived from the SIR model, with Kendall’s $\tau$ coefficient used as the evaluation metric. The two-stage model is trained on the BA synthetic network and evaluated on six real-world networks. The comparison results between ICAN and ICAN_w/o_Syner on different networks are presented in Fig. 2(b). Performance decline in ICAN_w/o_Syner is observed in most of the node ranking tasks. These results demonstrate that ICAN facilitates mutual reinforcement between causal representation learning and ranking prediction, resulting in more accurate ranking outcomes.

### v3

@@ 5.1 Description of datasets
We train the model on five synthetic graphs and evaluate it on eight real-world networks to validate its effectiveness.
Training datasets. Five representative synthetic network models are used for training, including Barabási–Albert (BA) [3], Extreme Homogeneous (EH) [45], Erdös-Rényi (ER) random-graph [13], Q-Snapback (QS) [44] and Random Hexagon (RH) [44]. The BA networks are generated using the preferential attachment mechanism, where new nodes are sequentially added and connected to existing nodes with a probability proportional to their current degrees, leading to a scale-free degree distribution. The EH networks are obtained by performing additional random edge rectifications on ER networks to transform their degree distribution from Poisson to near-uniform. The ER random graphs are generated by connecting each pair of nodes independently with a fixed probability $p$, producing networks with a binomial degree distribution. The QS networks consist of a directed backbone chain with multiple probabilistic ”snapback” edges linking newly added nodes to previously added ones within a defined range, thereby enhancing backward connectivity and robustness. The RH networks are composed of randomly connected hexagonal substructures that capture local clustering and spatial organization similar to lattice-like systems. All trained networks comprise 1,000 nodes with an average degree of 4.
@@
  • Karate [74]: A human social network, in which each node represents a member in the club and each link shows the relationship between two members.
  • Jazz [20]: A human collaboration network, in which each node represents a jazz musician and each link indicates that two musicians have performed together in the same band.
  • Email-univ [23]: An email communication network from the University Rovira i Virgili in Spain, where nodes represent users and edges indicate at least one email was exchanged.
  • USAir [37]: A directed weighted network represents the flight connections among 1,574 U.S. airports in 2010. Each node corresponds to an airport, and each directed edge indicates the existence of at least one flight from the origin to the destination airport. The edge weight quantifies the total number of flights operated on that route during the year.
  • Vidal [58]: A network represents an initial version of a proteomescale map of Human binary protein–protein interactions.
  • Email-dnc [57]: A directed unweighted network derived from the 2016 Democratic National Committee (DNC) email leak. Each node represents an individual user, and each directed edge indicates that at least one email was sent from one user to another.
  • Figeys [14]: A large-scale human protein–protein interaction network derived from mass spectrometry–based experimental assays, where nodes represent human proteins and edges denote experimentally detected physical interactions between them.
  • Oz [17]: A friendship network captures interpersonal relationships among 217 residents of a university residence hall, where each node represents an individual and each edge denotes a reported friendship connection between two residents.
@@ 5.2 Comparative methods
  • Degree Centrality (DC) [5]. It quantifies the influence of a node based on the number of direct connections it has to other nodes in the graph. The degree centrality of a node $u$ is formally defined as
@@
  • Betweenness Centrality (BC) [18]. It measures the importance of a node by quantifying how frequently it appears on the shortest paths between all pairs of nodes in the network.
@@
  • Eigenvector Centrality (EC) [6]. It assigns each node a centrality score based not only on its own connectivity but also on the importance of the nodes it is connected to. The EC score $\bm{X}_{u}$ of node $u$ satisfies
@@
  • $H$-index (HI) [27]. The $H$-index of a node considers the degrees of its neighbors and reflects how many of them are themselves influential. For a node $u$ with neighbors $j_{1},j_{2},\ldots,j_{k}$, the $H$-index is defined as
@@
  • $K$-Shell (KS) [34]. The $K$-shell algorithm operates through an iterative pruning process. It iteratively peels away network layers to assign each node a $k$-value indicating its coreness based on residual connections. A higher $K$-shell value indicates greater node influence.
  • GNN-Bet [47]. This method enables node ranking by employing a dual-path aggregation mechanism that separately propagates features along incoming and outgoing shortest paths using row-wise modified adjacency matrices, with final ranking scores generated through multiplicative combination of the aggregated path representations.
  • GNN-Close [47]. This approach achieves node ranking through a hierarchical feature aggregation scheme that combines normal adjacency operations in the initial layer with column-wise modified matrices in subsequent layers to constrain feature propagation along shortest paths, producing final ranking scores via summation of layer-wise outputs.
  • RCNN [72]. This method leverages convolutional neural networks to extract features from node-centric subgraph structures and ranks node importance based on learned representations.
  • CGNN [76]. This method enhances RCNN by incorporating additional GNN layers, improving the capture of topological dependencies for more accurate vital node identification.
  • AGNN [69]. This method combines a GCN-based autoencoder with a GNN ranking head to jointly learn structural embeddings and predict node importance using a listwise ranking loss.
  • STGCN [29]  [^1] Since AGNN and STGCN employ the adjacency matrix as the input to their autoencoder and do not explicitly address the issue of inconsistent numbers of nodes between training and testing networks, we instead incorporate the generated node features by the adjacency matrix as the input to the AGNN and STGCN models.. This method leverages a hybrid architecture combining Sparse Transformer layers for global dependency modeling and GCNs for local topological features to rank key nodes in complex networks.
@@ 5.3 Evaluation criterion
In this section, we introduce the evaluation metric used to measure the effectiveness of the predicted ranking results by ICAN, Kendall’s $\tau$ coefficient [35]. This non-parametric statistic is widely employed to measure the ordinal association between two ranking lists. It quantifies the similarity between two ranking orders by comparing the numbers of concordant and discordant pairs. The definition of Kendall’s $\tau$ is given by
@@
where $n$ is the number of items (or nodes), $N_{c}$ is the number of concordant pairs, $N_{d}$ is the number of discordant pairs, and $\tfrac{1}{2}n(n-1)$ is the total number of possible item pairs. For two rankings $\{(x_{i},y_{i})\}$ and $\{(x_{j},y_{j})\}$, a pair $(i,j)$ is considered concordant if $(x_{i}-x_{j})(y_{i}-y_{j})>0$, discordant if $(x_{i}-x_{j})(y_{i}-y_{j})<0$, and neither if $x_{i}=x_{j}$ or $y_{i}=y_{j}$. The value of Kendall’s $\tau$ ranges from 1-1 to $1$: a value of $1$ indicates complete agreement between rankings, 1-1 indicates complete disagreement, and 0 indicates no correlation. This metric is particularly useful for evaluating graph-based ranking models, as it reflects how well the predicted node ranking preserves the ground-truth ordering.
@@ 5.4 Implementation details
The parameter settings for ICAN and the comparative methods are summarized as follows.
For the ICAN method, set the number of hidden layers to be $l=5$ and each hidden layer to have $p=32$ neurons. Let
@@
The learning rate is set to $\mu=0.001$ and the maximum number of iterations is $T=10$. In addition, we set the recovery probability $\delta=1$ and the infection probability $\gamma=1.5\times\gamma_{c}$ to ensure effective spreading within the network. The infection threshold $\gamma_{c}$ is derived from mean-field theory [50] as
@@ 5.5.1 Training on BA-generated network and testing on various real-world networks
In the first experiment, we conduct experiments by training on the BA generated network and testing on different real-world networks to validate the model’s effectiveness. Table II presents the predicted Kendall’s correlation coefficients across different datasets. The experimental results comprehensively evaluate the proposed ICAN model against several baseline methods for node ranking across eight real-world networks. As shown in the table, ICAN demonstrates superior and consistent performance. Notably, ICAN achieves the highest Kendall’s coefficient on every individual dataset. This culminates in the best overall average Kendall’s coefficient of 0.7707, securing the top rank among all compared methods. The fact that ICAN, trained on a BA synthetic network, generalizes effectively to diverse real-world networks highlights its robust cross-network predictive capability. This can be attributed to the ICAN learning network-invariant node embeddings that transfer reliably across different networks.
Overall, in terms of average Kendall’s coefficient , ICAN ranks first in all datasets. From the experimental results above, we can summarize the following conclusions:
  • Compared to traditional topology-based methods, ICAN significantly outperforms heuristic baselines such as DC, BC and KS across all datasets. This demonstrates the limitations of methods that rely on single structural dimensions. In contrast, ICAN effectively captures complex, non-linear dependencies and higher-order topological information that traditional centrality measures fail to identify, proving the necessity of deep representation learning for node ranking.
  • Compared to deep learning-based methods, ICAN consistently outperforms state-of-the-art baselines, including RCNN, CGNN, AGNN, and STGCN. While frameworks like AGNN and STGCN also utilize autoencoders, they typically decouple representation learning from the ranking objective and lack explicit causal constraints. The superior performance of ICAN demonstrates that its influence-aware causal node embedding mechanism ensures the learned representations are causally linked to node importance, thereby enhancing robustness and generalization across diverse networks. Furthermore, by jointly minimizing causal reconstruction and ranking losses, ICAN adopts a unified optimization strategy that fosters mutual reinforcement between feature extraction and ranking prediction. This alignment ensures that the network-invariant causal representations are directly tailored for the downstream ranking task, leading to superior cross-network transferability.
@@
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Figeys Oz Average Rank
DC 0.6749 0.7903 0.6236 0.6119 0.4648 0.5613 0.5950 0.7587 0.6351 6
BC 0.5685 0.4643 0.2374 0.4407 0.5876 0.4032 0.5321 0.5060 0.4675 12
EC 0.8069 0.8212 0.2534 0.6554 0.2452 0.5503 0.6546 0.7684 0.5944 7
HI 0.6654 0.8271 0.6854 0.6284 0.7568 0.5672 0.6245 0.6299 0.6946 4
KS 0.6421 0.7713 0.6828 0.6369 0.7619 0.4988 0.6299 0.7069 0.6663 5
GNN-Bet 0.6530 0.6789 0.3645 0.5350 0.2866 0.3662 0.5620 0.4902 0.4921 10
GNN-Close 0.5352 0.6551 0.3629 0.5818 0.2429 0.4442 0.5453 0.3832 0.4688 11
RCNN 0.7528 0.8455 0.7522 0.6709 0.9035 0.5432 0.6025 0.7914 0.7328 2
CGNN 0.7058 0.7911 0.6488 0.5835 0.9442 0.5649 0.5959 0.7322 0.6958 3
AGNN 0.6505 0.7584 0.3100 0.6429 0.4108 0.4946 0.5989 0.2744 0.5176 9
STGCN 0.5640 0.7000 0.7327 0.3972 0.5784 0.4248 0.5975 0.6504 0.5806 8
ICAN 0.8090 0.8504 0.7625 0.6758 0.9524 0.5740 0.6708 0.8051 0.7625 1
@@ 5.5.2 Training on ER-generated network and testing on various real-world networks
Next, we conduct experiments by training on the ER generated network and testing on different real-world networks. Table III presents the Kendall’s $\tau$ coefficients of various methods evaluated across eight real-world networks. As shown, ICAN attains the highest Kendall’s $\tau$ values on four out of eight datasets, with an average coefficient of 0.7732, ranking first overall. This clearly demonstrates its strong generalization ability when trained on the ER-generated network and tested on structurally diverse real-world networks.
@@
Methods Karate Jazz Email-univ USAir Vidal Email-dnc Figeys Oz Average Rank
DC 0.6749 0.7903 0.6236 0.6119 0.4648 0.5613 0.5950 0.7587 0.6351 6
BC 0.5685 0.4643 0.2374 0.4407 0.5876 0.4032 0.5321 0.5060 0.4675 11
EC 0.8069 0.8212 0.2534 0.6554 0.2452 0.5503 0.6546 0.7684 0.5944 7
HI 0.6654 0.8271 0.6854 0.6284 0.7568 0.5672 0.6245 0.8017 0.6946 3
KS 0.6421 0.7713 0.6828 0.6369 0.7619 0.4988 0.6299 0.7069 0.6663 5
GNN-Bet 0.6137 0.6394 0.3744 0.5565 0.4125 0.4566 0.5365 0.5691 0.5198 10
GNN-Close 0.6887 0.2779 0.4550 0.5470 0.4079 0.4122 0.5343 0.1673 0.4363 12
RCNN 0.7094 0.8014 0.8141 0.6823 0.9119 0.5490 0.5576 0.7507 0.7221 2
CGNN 0.6749 0.6214 0.6323 0.5835 0.9442 0.5649 0.5959 0.7322 0.6687 4
AGNN 0.7622 0.4889 0.2950 0.6829 0.4008 0.4534 0.5091 0.6549 0.5309 9
STGCN 0.5568 0.6525 0.7141 0.3952 0.5792 0.3735 0.3833 0.6751 0.5412 8
ICAN 0.8092 0.8460 0.8228 0.6545 0.9524 0.5540 0.6312 0.7820 0.7565 1
@@ 5.6.2 Ablation study for the feature–task co-optimization mechanism
In this study, we integrate the causal ranking loss and causal reconstruction loss within a unified framework to enable causal feature extraction and ranking tasks. To evaluate the effectiveness of the proposed end-to-end feature–task co-optimization mechanism, we conduct an ablation study by transforming the integrated model into a two-stage implementation, in which the feature extraction and ranking processes are decoupled, thereby removing the synergistic optimization effect. This variant is referred to as ICAN_w/o_Syner. In the first stage, the causal reconstruction loss and a regularization term are employed as the loss function to optimize the causal node embedding module, yielding low-dimensional node representations that comprehensively encapsulate the topological properties of the network. During the second stage, the learned node embeddings are fed into the ranking prediction module, which is optimized using the causal ranking loss along with a regularization term to generate the predicted node influence score. These predictions are then compared against the node influence score derived from the SIR model, with Kendall’s $\tau$ coefficient used as the evaluation metric. The two-stage model is trained on the BA synthetic network and evaluated on eight real-world networks. The comparison results between ICAN and ICAN_w/o_Syner on different networks are presented in Fig. 2(b). Performance decline in ICAN_w/o_Syner is observed in most of the node ranking tasks. These results demonstrate that ICAN facilitates mutual reinforcement between causal representation learning and ranking prediction, resulting in more accurate ranking outcomes.